# Test-Time Compute: from System-1 Thinking to System-2 Thinking

### v2

@@ 5.1 Generalizable System-2 Model
Currently, most o1-like models exhibit strong reasoning abilities only in specific domains such as math and code and struggle to adapt to cross-domain or general tasks. The key to addressing this issue lies in enhancing the generalization ability of verifiers or critics (LeVine et al., 2023; Kim et al., 2024d; Chen et al., 2024c). Currently, some works utilize multi-objective training (Wang et al., 2024c), model ensemble (Lin et al., 2024b) or regularization constraints (Yang et al., 2024a; Jia, 2024) to make verifiers more generalizable. Nevertheless, there is still significant room for improvement in the generalization of the verifier. Additionally, weak-to-strong generalization (Burns et al., 2023) is a topic worth further exploration. People are no longer satisfied with solving mathematical problems with standard answers; they hope System-2 models can assist in scientific discovery and the proofs of mathematical conjectures. In such cases, even human experts struggle to provide accurate feedback, while weak-to-strong generalization offers a promising direction to address this issue (Tang et al., 2025). We think that more generalized System-2 models may not rely on a single feedback source but instead obtain multi-source feedback through interactions between LLM-based agents and tools, experts, or other agents (Nathani et al., 2023; Lan et al., 2024).
5.2 Multimodal Reasoning
In System-1 thinking, TTA has been successfully applied to multimodal LLMs, improving performance in tasks such as zero-shot image classification, image-text retrieval, and image captioning (Zhao et al., 2024a). However, test-time compute methods in System-2 thinking remain limited to text modalities. Visual, speech, and other modalities are crucial for model understanding and interaction with the world. To achieve cognitive intelligence, System-2 models must be able to fully integrate multimodal information for reasoning. The exploration of multimodal CoT (Zhang et al., 2024l; Wu et al., 2024b; Mondal et al., 2024; Lee et al., 2024; Gao et al., 2024b) and multimodal critics or verifiers (Xiong et al., 2024) open up the possibility of building multimodal System-2 models. Xu et al. (2024b) are the first to apply test-time compute to visual reasoning tasks. They divide the visual reasoning process into four stages: task summary, caption, reasoning, and answer conclusion. They propose a stage-level beam search method, which repeatedly samples at each stage and selects the best result for the next stage. Nowadays, Qwen team has released the open-weight multimodal reasoning model QVQ (Qwen, 2024), OpenAI and Kimi (Team et al., 2025) have released their multimodal reasoning products. We believe test-time compute still holds significant potential for development in multimodal reasoning. For example, incorporating more modalities like speech and video into reasoning tasks, applying successful methods such as reflection mechanisms and tree search (Yao et al., 2024; Dong et al., 2024a) to multimodal reasoning, or aligning the multimodal reasoning process with human cognitive processes. Besides understanding and reasoning tasks, Xie et al. (2025) and Guo et al. (2025b) show test-time compute can improve image generation performance, with great potential for multimodal generation in the future.
5.3 Efficiency and Performance Trade-off
@@
The successful application of test-time compute shows that sacrificing reasoning efficiency can lead to better reasoning performance. However, researchers continue to seek a balance between performance and efficiency, aiming to achieve optimal performance under a fixed reasoning latency budget. This requires adaptively allocating computational resources for each sample. Damani et al. (2024) train a lightweight module to predict the difficulty of a question, and allocate computational resources according to its difficulty. Zhang et al. (2024f) further extend the allocation targets to more hyperparameters. Chen et al. (2025b) and Wang et al. (2025) systematically evaluate the over-thinking and under-thinking phenomena in o1-like models, where the former leads models to overcomplicate simple problems, and the latter causes frequent switching of reasoning paths on difficult problems, thereby reducing reasoning efficiency. O1-Pruner (Luo et al., 2025b) propose the length-penalty PPO loss to shorten reasoning processes while maintaining accuracy. There are still many open questions worth exploring, such as how to integrate inference acceleration strategies, e.g. model compression (Li et al., 2024f; Huang et al., 2024c; Li et al., 2025b), token pruning (Fu et al., 2024; Zhang et al., 2024d), and speculative decoding (Leviathan et al., 2023; Xia et al., 2024) with test-time compute, and how to allocate optimal reasoning budget according to problem difficulty.
5.4 Scaling Law
@@
Unlike training-time computation scaling, test-time compute still lacks a universal scaling law. Some works have attempted to derive scaling laws for specific test-time compute strategies (Wu et al., 2024c; Levi, 2024). Brown et al. (2024) demonstrate that the performance has an approximately log-linear relationship with repeated sampling times. Chen et al. (2024e) models repeated sampling as a knockout tournament and league-style algorithm, proving theoretically that the failure probability of repeated sampling follows a power-law scaling. Snell et al. (2024) investigate the scaling laws of repeated sampling and self-correction, and propose the computing-optimal scaling strategy. There are two major challenges to achieving a universal scaling law: first, current test-time compute strategies are various, each with different mechanisms to steer the model; thus, it lacks a universal framework for describing them; second, the performance of test-time compute is affected by a variety of factors, including the difficulty of samples, the accuracy of feedback signals, and decoding hyperparameters, and we need empirical studies to filter out the critical factors.
5.5 Strategy Combination
@@
Different test-time compute strategies are suited to various tasks and scenarios, so combining multiple strategies is one way to achieve better System-2 thinking. For example, Marco-o1 (Zhao et al., 2024b) combines the MCTS and self-correction, using MCTS to plan reasoning processes, and self-correction to improve the accuracy of each step. TPO (Li et al., 2025a) combines BoN sampling and self-correction. Moreover, test-time adaptation strategies in System-1 models can also be combined with test-time reasoning strategies. Akyürek et al. (2024) combine test-time training with repeated sampling. They further optimize the language modeling loss on test samples, then generate multiple candidate answers through data augmentation, and finally determine the answer by majority voting. They demonstrate the potential of test-time training in reasoning tasks, surpassing the human average on the ARC challenge. Therefore, we think that for LLM reasoning, it is crucial to focus not only on emerging test-time strategies but also on test-time adaptation methods. By effectively combining these strategies, we can develop System-2 models that achieve or surpass o1-level performance.
6 Benchmarks and Open-source Frameworks
6.1 Benchmarks
@@
In System-1 models, distribution shifts include adversarial robustness, cross-domain and cross-lingual scenarios. In the field of CV, ImageNet-C (Hendrycks and Dietterich, 2019), ImageNet-R (Hendrycks et al., 2021a), ImageNet-Sketch (Wang et al., 2019) are common datasets for TTA. Yu et al. (2023) propose a benchmark to conduct a unified evaluation of TTA methods across different TTA settings and backbones on 5 image classification datasets. For NLP tasks, TTA is primarily applied in QA and machine translation tasks, with commonly used datasets such as MLQA (Lewis et al., 2020), XQuAD (Artetxe et al., 2020), MRQA (Fisch et al., 2019), CCMatrix (Schwenk et al., 2021) and Ted Talks (Qi et al., 2018).
@@
RewardBench (Lambert et al., 2024) collects 20.2k prompt-choice-rejection triplets covering tasks such as dialogue, reasoning, and safety. It evaluates the accuracy of reward models in distinguishing between chosen and rejected responses. RM-Bench (Liu et al., 2024d) further evaluates the impact of response style on reward models. RMB (Zhou et al., 2024) extends the evaluation to the more practical BoN setting, where reward models are required to select the best response from multiple candidates. CriticBench (Lin et al., 2024c) is specifically designed to evaluate a critic model’s generation, critique, and correction capabilities. For PRM, Song et al. (2025a) propose PRMBench, which evaluates PRMs whether they can identify the earliest incorrect reasoning step in math tasks. ProcessBench (Zheng et al., 2024) provides a more fine-grained evaluation, including redundancy, soundness, and sensitivity. In addition, there are benchmarks for evaluating multimodal feedback modeling, such as VL-RewardBench (Li et al., 2024d) and MJ-Bench (Chen et al., 2024f).
@@
Reasoning capability is the core of System-2 models, including mathematics, code, commonsense, planning, etc (Zeng et al., 2024). Math reasoning is one of the most compelling reasoning tasks. With the advancements in LLM and test-time compute, the accuracy on some previously challenging benchmarks, like GSM8K (Cobbe et al., 2021) and MATH (Hendrycks et al., 2021b), have surpassed the 90% mark. Thus, more difficult college admissions exam (Zhang et al., 2023b; Arora et al., 2023; Azerbayev et al., 2024) and competition-level (Gao et al., 2024a) math benchmarks have been proposed. Some competition-level benchmarks are not limited to textual modalities in algebra, logic reasoning, and word problems. For instance, OlympiadBench (He et al., 2024), OlympicArena (Huang et al., 2024b) and AIME (Zamil and Rabby, 2024) provide images for geometry problems, incorporating visual information to aid in problem-solving, while AlphaGeometry (Trinh et al., 2024) employs symbolic rules for geometric proofs. The most challenging benchmark currently is FrontierMath (Glazer et al., 2024), with problems crafted by mathematicians and covering major branches of modern mathematics. Even the most advanced o3 has not achieved 30% accuracy.
@@
Code ability is a key aspect of LLM reasoning, with high practical value, covering code completion (Ding et al., 2023; Zhang et al., 2023a; Gong et al., 2024a), code reasoning (Gu et al., 2024), and code generation (Chen et al., 2021; Austin et al., 2021) tasks. Among these, code generation gains more attention. HumanEval (Chen et al., 2021) and MBPP (Austin et al., 2021) provide natural language descriptions of programming problems, requiring LLMs to generate corresponding Python code and use unit tests for evaluation. MultiPL-E (Cassano et al., 2022) extend them to 18 program languages. EvalPlus (Liu et al., 2024a) automatically augments test cases to assess the robustness of the generated code. Recently, some studies collect benchmarks from open-source projects, which are closed to realistic applications and more challenging due to complex function calls, such as DS-1000 (Lai et al., 2023), CoderEval (Yu et al., 2024a), EvoCodeBench (Li et al., 2024a) and BigCodeBench (Zhuo et al., 2025).
@@
Commonsense reasoning requires LLMs to possess both commonsense knowledge and reasoning abilities. Early benchmarks (Zellers et al., 2019; Talmor et al., 2019; Sakaguchi et al., 2021; Bisk et al., 2020) focus on evaluating LLMs’ commonsense ability. StrategyQA (Geva et al., 2021) collects more complex and subtle multi-hop reasoning questions. MMLU (Hendrycks et al., 2021b) and MMLU-Pro (Wang et al., 2024n) cover commonsense reasoning questions across various domains, including STEM, the humanities, the social sciences, etc. Planning aims to enable LLMs to take optimal actions based on the current state and environment to complete tasks. Current planning benchmarks primarily focus on synthetic tasks, such as Blocksworld (Valmeekam et al., 2023b), Crosswords, and Game-of-24 (Yao et al., 2023).
6.2 Projects
OpenR
(Wang et al., 2024d) [^2] https://github.com/openreasoner/openr is an open-source test-time reasoning framework that integrates various test-time compute strategies, PRM training, and improvement training. It currently supports beam search, BoN, MCTS, and rStar, and implements popular online reinforcement learning algorithms like APPO, GRPO, and TPPO.
RLHFlow
(Dong et al., 2024b) offers a comprehensive framework for reward modeling [^3] https://github.com/RLHFlow/RLHF-Reward-Modeling and online RLHF training [^4] https://github.com/RLHFlow/Online-RLHF. Its standout feature is the integration of various reward model training methods, including the vanilla preference reward model, multi-objective reward models, PRM, etc.
OpenRLHF
(Hu et al., 2024a) [^5] https://github.com/OpenRLHF/OpenRLHF also integrates reward modeling and RLHF training but focuses more on the efficient implementation of reinforcement learning algorithms and training tricks. Its strength lies in the integration of distributed training and efficient fine-tuning, enabling users to easily train large language models with more than 70B parameters.
7 Conclusion

### v3

@@ 5.1. Generalizable System-2 Model
Additionally, weak-to-strong generalization (Burns et al., 2023) is a topic worth further exploration. People are no longer satisfied with solving mathematical problems with standard answers; they hope System-2 models can assist in scientific discovery and the proofs of mathematical conjectures. In such cases, even human experts struggle to provide accurate feedback, while weak-to-strong generalization offers a promising direction to address this issue (Tang et al., 2025).
5.2. Multimodal Test-time Compute
Visual, audio, video and other modalities are crucial for models understanding and interaction with the world. Like LLMs, large multimodal models (LMMs) also face challenges in scaling training compute. Thus, enabling efficient test-time scaling for multimodal tasks has become an advanced research topic. In System-1 thinking, TTA has been successfully applied to LMMs, improving performance in tasks such as zero-shot image classification, image-text retrieval, and image captioning (Zhao et al., 2024a). For multimodal reasoning tasks, the exploration of multimodal CoT (Zhang et al., 2024j; Wu et al., 2024b; Mondal et al., 2024; Lee et al., 2024; Gao et al., 2024b; Zhang et al., 2024k) and multimodal critics or verifiers (Xiong et al., 2024; Tu et al., 2025) open up the possibility of building multimodal System-2 models. For example, IoT (Zhou et al., 2024b) and SketchPAD (Hu et al., 2024a) utilize visual tools to obtain critical visual information and integrate it into CoT to refine multimodal reasoning. VisualPRM (Wang et al., 2025a) constructs process annotation data through Monte Carlo estimation and trains a lightweight PRM verifier; and cutting-edge LMMs such as GPT-4o and Qwen-VL can serve as critics. Building on these foundations, repeated sampling (Lin et al., 2025) and tree search (Yao et al., 2024; Dong et al., 2024b) can effectively further improve LMMs’ multimodal reasoning performance. Xu et al. (2024a) divide the visual reasoning process into four stages: task summary, caption, reasoning, and answer conclusion. They propose a stage-level beam search method, which repeatedly samples at each stage and selects the best result for the next stage. Audio-Reasoner (Xie et al., 2025b) transfers the four-stage CoT framework to the audio modality. To equip LMMs with the self-correction ability, Deepseek-R1’s training recipe has also been practiced in LMMs (Ma et al., 2025a; Chen et al., 2025d; Shen et al., 2025a). OpenVLThinker (Deng et al., 2025) and Vision-R1 (Huang et al., 2025b) convert images into text captions, enabling text-only LRMs to incorporate visual information for generating CoTs. This serves as cold-start data for further reinforcement learning to train multimodal reasoning models. LMM-R1 (Peng et al., 2025b) demonstrates that the reasoning ability acquired through text-only reinforcement learning can serve as an effective cold start for multimodal reinforcement learning training.
Besides understanding and reasoning tasks, multimodal generation tasks can also benefit from test-time compute. Mainstream multimodal generative models are divided into diffusion-based (Peebles and Xie, 2023; Yang et al., 2025c) and autoregressive-based (Tian et al., 2024; Xie et al., 2025c) models. While diffusion models lack explicit CoTs, they can expand the search space at test time by sampling various Gaussian noises (Xie et al., 2025a). For video generation, due to the high computational cost of repeatedly sampling full videos, Cong et al. (2025) use PRM to verify frames individually and apply beam search to prune low-scoring frames. Liu et al. (2025d) propose frame tree search, using critics to evaluate videos at early, middle, and late stages with different criteria, retaining only the top-k scoring frames at each step. Autoregressive models treat intermediate generated images as reasoning steps. Guo et al. (2025b) propose a Potential Assessment Reward Model (PARM) to evaluate the quality potential of these intermediate images and introduce a reflection mechanism that enables the generative model to self-correct low-quality images.
5.3. Efficient Test-time Compute
@@
The successful application of test-time compute shows that sacrificing reasoning efficiency can lead to better reasoning performance. However, researchers continue to seek a balance between performance and efficiency, aiming to achieve optimal performance under a fixed reasoning latency budget. This requires adaptively allocating computational resources for each sample. Damani et al. (2024) train a lightweight module to predict the difficulty of a question, and allocate computational resources according to its difficulty. Zhang et al. (2024l) further extend the allocation targets to more hyperparameters. Chen et al. (2025e) and Wang et al. (2025b) systematically evaluate the over-thinking and under-thinking phenomena in LRMs, where the former leads models to overcomplicate simple problems, and the latter causes frequent switching of reasoning paths on difficult problems, thereby reducing reasoning efficiency. These phenomena even cause LRM to perform poorly when dealing with simple problems (Zhang et al., 2025b). There are still many open questions worth exploring, such as how to integrate inference acceleration strategies, e.g. model compression (Li et al., 2024c; Huang et al., 2024c; Li et al., 2025f), token pruning (Fu et al., 2024; Zhang et al., 2024i), and speculative decoding (Leviathan et al., 2023; Xia et al., 2024) with test-time compute, and how to allocate optimal reasoning budget according to problem difficulty (Wang et al., 2024b; Han et al., 2024; Cheng and Durme, 2024).
Long CoT leads to inference latency and memory footprints of key-value cache. Recently, numerous studies have explored various strategies to reduce the reasoning length. Early work primarily focus on SFT models (Yu et al., 2024c; Liu et al., 2024a; Kang et al., 2025a; Xia et al., 2025; Ma et al., 2025c; Munkhbat et al., 2025; Cui et al., 2025; Yang et al., 2025a) using variable-length CoT data. In contrast, mainstream approaches incorporate length-based rewards into reinforcement learning (RL) (Team et al., 2025; Luo et al., 2025c; Arora and Zanette, 2025; Yeo et al., 2025; Yi et al., 2025; Yuan et al., 2025c) to encourage concise responses. Xiao et al. (2025) measure the complexity of vision-language reasoning tasks and design a complexity-aware curriculum learning and GRPO algorithm. They organize training samples from hard to easy and dynamically adjust the reward function and KL term weights based on task complexity. L1 (Aggarwal and Welleck, 2025) and Elastic Reasoning (Xu et al., 2025a) can enable precise control over response length based on a given token budget. Further, SelfBudgeter (Li et al., 2025c) can autonomously estimate the optimal token budget before generation, and allow users to choose whether to wait for the full response or terminate early during the generation process. However, both SFT and RL require substantial computational resources. Several studies have instead explored training-free approaches for efficient reasoning, including prompting (Han et al., 2024; Xu et al., 2025d; Aytes et al., 2025; Lee et al., 2025), model merging (Wu et al., 2025; Team et al., 2025) and early exit (Fu et al., 2025; Ma et al., 2025b; Yang et al., 2025b). Notably, DEER (Yang et al., 2025b) makes early-exit decisions during CoT generation based on the confidence of intermediate answers, effectively reducing reasoning overhead. Subsequently,  Dai et al. (2025) introduces a new RL paradigm that encourages the model to incentivize early thinking termination when appropriate. Another promising approach to achieving efficiency is hybrid reasoning, which dynamically switches between concise responses and long-chain reasoning based on task complexity. These methods (Lou et al., 2025; Luo et al., 2025b; Jiang et al., 2025; Fang et al., 2025; Zhang et al., 2025a) follow a similar pipeline: they start with SFT on a mixed dataset of long and short CoT samples as a cold start, and then refine the policy optimization algorithm to train the model to acquire hybrid reasoning capabilities.
6. Future Directions
6.1. Test-time Scaling Law
@@
Unlike training-time computation scaling, test-time compute still lacks a universal scaling law. Some works have attempted to derive scaling laws for specific test-time compute strategies (Wu et al., 2024c; Levi, 2024). Brown et al. (2024) demonstrate that the performance has an approximately log-linear relationship with repeated sampling times. Chen et al. (2024d) models repeated sampling as a knockout tournament and league-style algorithm, proving theoretically that the failure probability of repeated sampling follows a power-law scaling. Snell et al. (2024) investigate the scaling laws of repeated sampling and self-correction, and propose the computing-optimal scaling strategy. There are two major challenges to achieving a universal scaling law: first, current test-time compute strategies are various, each with different mechanisms to steer the model; thus, it lacks a universal framework for describing them; second, the performance of test-time compute is affected by a variety of factors, including the difficulty of samples, the accuracy of feedback signals, and decoding hyperparameters, and we need empirical studies to filter out the critical factors.
6.2. Strategy Combination
@@
Different test-time compute strategies are suited to various tasks and scenarios, so combining multiple strategies is one way to achieve better System-2 thinking. For example, Marco-o1 (Zhao et al., 2024b) combines the MCTS and self-correction, using MCTS to plan reasoning processes, and self-correction to improve the accuracy of each step. TPO (Li et al., 2025d) combines BoN sampling and self-correction. Moreover, test-time adaptation strategies in System-1 models can also be combined with test-time reasoning strategies. Akyürek et al. (2024) combine test-time training with repeated sampling. They further optimize the language modeling loss on test samples, then generate multiple candidate answers through data augmentation, and finally determine the answer by majority voting. They demonstrate the potential of test-time training in reasoning tasks, surpassing the human average on the ARC challenge. Therefore, we think that for LLM reasoning, it is crucial to focus not only on emerging test-time strategies but also on test-time adaptation methods. By effectively combining these strategies, we can develop System-2 models that achieve or surpass o1-level performance.
6.3. New Test-time Compute Paradigms
Latent CoT reasoning has emerged as a promising paradigm for test-time compute. By performing internal reasoning in latent space, it overcomes the limitations of conventional explicit CoT in terms of computational efficiency and abstract reasoning capability (Goyal et al., 2024; Xu et al., 2025b; Shen et al., 2025b; Xu et al., 2025c). Coconut (Hao et al., 2024) performs reasoning in a continuous latent space by directly feeding the model’s final hidden state as the next-step input embedding, thereby giving rise to advanced reasoning patterns. CCoT (Cheng and Durme, 2024) introduces a framework that generates continuous and variable-length contemplation tokens as compressed representations of reasoning chains, enabling efficient and accurate reasoning. LightThinker (Zhang et al., 2025d) improves LLM efficiency by compressing intermediate reasoning steps into compact gist tokens via specialized data construction and attention mask design, enabling the model to discard verbose chains while maintaining the ability to reason over condensed representations. Latent CoT currently faces several challenges, including the difficulty of direct supervision, limited generalization, and concerns over interpretability (Chen et al., 2025f). This calls for further exploration of token-level strategies and internal mechanisms to enable more robust and transparent latent reasoning.
7. Benchmarks
@@
In System-1 models, distribution shifts include adversarial robustness, cross-domain and cross-lingual scenarios. In the field of CV, ImageNet-C (Hendrycks and Dietterich, 2019), ImageNet-R (Hendrycks et al., 2021a), ImageNet-Sketch (Wang et al., 2019) are common datasets for TTA. Yu et al. (2023) propose a benchmark to conduct a unified evaluation of TTA methods across different TTA settings and backbones on 5 image classification datasets. For NLP tasks, TTA is primarily applied in QA and machine translation tasks, with commonly used datasets such as MLQA (Lewis et al., 2020), XQuAD (Artetxe et al., 2020), MRQA (Fisch et al., 2019) and CCMatrix (Schwenk et al., 2021).
@@
RewardBench (Lambert et al., 2024) collects 20.2k prompt-choice-rejection triplets covering tasks such as dialogue, reasoning, and safety. It evaluates the accuracy of reward models in distinguishing between chosen and rejected responses. RM-Bench (Liu et al., 2024d) further evaluates the impact of response style on reward models. RMB (Zhou et al., 2024a) extends the evaluation to the more practical BoN setting, where reward models are required to select the best response from multiple candidates. CriticBench (Lin et al., 2024b) is specifically designed to evaluate a critic model’s generation, critique, and correction capabilities. For PRM, Song et al. (2025a) propose PRMBench, which evaluates PRMs whether they can identify the earliest incorrect reasoning step in math tasks. ProcessBench (Zheng et al., 2024) provides a more fine-grained evaluation, including redundancy, soundness, and sensitivity. In addition, there are benchmarks for evaluating multimodal feedback modeling, such as VL-RewardBench (Li et al., 2024g), VCR-Bench (Qi et al., 2025) and MJ-Bench (Chen et al., 2024a).
@@
Reasoning capability is the core of System-2 models, including mathematics, code, commonsense, planning, etc (Zeng et al., 2024). Math reasoning is one of the most compelling reasoning tasks. With the advancements in LLM and test-time compute, the accuracy on some previously challenging benchmarks, like GSM8K (Cobbe et al., 2021) and MATH (Hendrycks et al., 2021b), have surpassed the 90% mark. Thus, more difficult college admissions exam (Zhang et al., 2023b; Arora et al., 2023; Azerbayev et al., 2024) and competition-level (Gao et al., 2024d) math benchmarks have been proposed. Some competition-level benchmarks are not limited to textual modalities in algebra, logic reasoning, and word problems. For instance, OlympiadBench (He et al., 2024), OlympicArena (Huang et al., 2024b) and AIME (Zamil and Rabby, 2024) provide images for geometry problems, incorporating visual information to aid in problem-solving, while AlphaGeometry (Trinh et al., 2024) employs symbolic rules for geometric proofs. The most challenging benchmark currently is FrontierMath (Glazer et al., 2024), with problems crafted by mathematicians and covering major branches of modern mathematics. Even the most advanced o3 has not achieved 30% accuracy.
@@
Code ability is a key aspect of LLM reasoning, with high practical value, covering code completion (Ding et al., 2023; Zhang et al., 2023a; Gong et al., 2024b), code reasoning (Gu et al., 2024; Wei et al., 2025), and code generation (Chen et al., 2021; Austin et al., 2021) tasks. Among these, code generation gains more attention. HumanEval (Chen et al., 2021) and MBPP (Austin et al., 2021) provide natural language descriptions of programming problems, requiring LLMs to generate corresponding Python code and use unit tests for evaluation. MultiPL-E (Cassano et al., 2022) extend them to 18 program languages. EvalPlus (Liu et al., 2024c) automatically augments test cases to assess the robustness of the generated code. Recently, some studies collect benchmarks from open-source projects, which are closed to realistic applications and more challenging due to complex function calls, such as DS-1000 (Lai et al., 2023), CoderEval (Yu et al., 2024b), EvoCodeBench (Li et al., 2024d) and BigCodeBench (Zhuo et al., 2025).
@@
Commonsense reasoning requires LLMs to possess both commonsense and reasoning abilities. StrategyQA (Geva et al., 2021) collects complex and subtle multi-hop reasoning questions. MMLU (Hendrycks et al., 2021b) and MMLU-Pro (Wang et al., 2024j) cover commonsense reasoning questions across various domains, including STEM, the humanities, the social sciences, etc. Planning aims to enable LLMs to take optimal actions based on the current state and environment to complete tasks. Current planning benchmarks primarily focus on synthetic tasks, such as Blocksworld (Valmeekam et al., 2023b), Crosswords, and Game-of-24 (Yao et al., 2023).
8. Conclusion